OpenStudy (anonymous):

(4 to the -1 over 3)-2

OpenStudy (anonymous):

so\[\frac{ 4^{-1} }{ 3^{-2 }}\]

OpenStudy (anonymous):

the rule of exponents is if it is a negative, it flips to the reciprocal. then 4^-1 goes to 1/4^1, numerators go to denominators and visaversa

OpenStudy (anonymous):

\[\frac{ 4^{-1} }{ 3 }-2\] = \[\frac{ 1 }{ 4*3 }-2\]

OpenStudy (anonymous):

4 to the -1 over three and in parentheses and all to the -2

OpenStudy (anonymous):

yeah if that is what's asked then that is right.

OpenStudy (anonymous):

but the rule still applies

OpenStudy (anonymous):

\[(\frac{ 4^{-1} }{ 3 })^{-2}\]=\[(\frac{ 1 }{ 4*3 })^{-2}\]=\[(\frac{ 1 }{ 12 })^{-2}\]=\[(\frac{ 12 }{ 1 })^{2}\]=144

OpenStudy (anonymous):

fmadson, i would learn to type the equation in numbers how it is to look, it would help everyone involved

OpenStudy (anonymous):

ok thanks